Run Log

From SBS wiki
Jump to: navigation, search

Return to main SBS wiki.

Production data

Carbon and Dummy Runs

Run Info Beam Settings DAQ BigBite Spectrometer Settings Super BigBite Spectrometer Settings Notes
Num Date Time Target Ebeam (GeV) Ibeam (uA) CODA Config Pre- scales Angle (º) Distance (m) Current (A) B-Field (T) Threshold (mV) SBS Angle (º) SBS Dist (m) Current (A) B-Field (T) HCal Angle (º) HCal Dist (m) Threshold (mV) Notes
11445 10/21/21 09:42-10:17 C 0.5% 3.728 6 uA GMN1 ps1=0, ps2-8=-1 36.0 1.80 750 -500 31.9 2.25 630 31.9 11 -20

LH2 Runs

Run Info Beam Settings DAQ BigBite Spectrometer Settings Super BigBite Spectrometer Settings Notes
Num Date Time Target Ebeam (GeV) Ibeam (uA) CODA Config Pre- scales Angle (º) Distance (m) Current (A) B-Field (T) Threshold (mV) SBS Angle (º) SBS Dist (m) Current (A) B-Field (T) HCal Angle (º) HCal Dist (m) Threshold (mV) Notes
11338 10/19/21 00:14-06:15 Out of Beam 0 0 GMN1 ps1=0, ps2-8=-1 51.0 1.85 750 -275 33.5 2.25 0 34.5 13.5 -20 run going while ops establishes beam to the hall
11340 10/19/21 06:16-13:52 Out of Beam 0 0 GMN1 ps1=0, ps2-8=-1 51.0 1.85 750 -275 33.5 2.25 0 34.5 13.5 -20 2-pass setup. Beam rarely off and on while establishing
11343 10/19/21 13:09-13:20 Out of Beam 0 0 GMN1 ps1=0, ps2-8=-1 51.0 1.85 750 -275 33.5 2.25 0 34.5 13.5 -20 2-pass setup. run while ops did calibrations
11420 10/21/21 04:30 LH2 3.728 1 uA GMN1 ps1=0, ps2-8=-1 36.0 1.80 750 -275 31.9 2.25 630 31.9 11 -20
11421 10/21/21 04:49 LH2 3.728 2 uA GMN1 ps1=0, ps2-8=-1 36.0 1.80 750 -450 31.9 2.25 630 31.9 11 -20
11422 10/21/21 05:00 LH2 3.728 3 uA GMN1 ps1=0, ps2-8=-1 36.0 1.80 750 -450 31.9 2.25 630 31.9 11 -20
11423 10/21/21 05:11 LH2 3.728 4 uA GMN1 ps1=0, ps2-8=-1 36.0 1.80 750 -450 31.9 2.25 630 31.9 11 -20
11425 10/21/21 05:37 LH2 3.728 5 uA GMN1 ps1=0, ps2-8=-1 36.0 1.80 750 -500 31.9 2.25 630 31.9 11 -20
11427 10/21/21 05:49-05:51 LH2 3.728 5 uA GMN1 ps1=0, ps2-8=-1 36.0 1.80 750 -500 31.9 2.25 630 31.9 11 -20
11428 10/21/21 05:55-05:58 LH2 3.728 5 uA GMN1 ps1=0, ps2-8=-1 36.0 1.80 750 -500 31.9 2.25 630 31.9 11 -20
11429 10/21/21 05:59-06:02 LH2 3.728 5 uA GMN1 ps1=0, ps2-8=-1 36.0 1.80 750 -500 31.9 2.25 630 31.9 11 -20
11430 10/21/21 06:05-06:08 LH2 3.728 6.5 uA GMN1 ps1=0, ps2-8=-1 36.0 1.80 750 -500 31.9 2.25 630 31.9 11 -20
11431 10/21/21 06:15-06:27 LH2 3.728 6.5 uA GMN1 ps1=0, ps2-8=-1 36.0 1.80 750 -500 31.9 2.25 630 31.9 11 -20
11432 10/21/21 06:31-06:50 LH2 3.728 6.5 uA GMN1 ps1=0, ps2-8=-1 36.0 1.80 750 -500 31.9 2.25 630 31.9 11 -20
11433 10/21/21 06:51-06:53 LH2 3.728 6.5 uA GMN1 ps1=0, ps2-8=-1 36.0 1.80 750 -500 31.9 2.25 630 31.9 11 -20
11434 10/21/21 06:57-07:01 LH2 3.728 6.5 uA GMN1 ps1=0, ps2-8=-1 36.0 1.80 750 -500 31.9 2.25 630 31.9 11 -20
11435 10/21/21 07:03-07:08 LH2 3.728 6.5 uA GMN1 ps1=0, ps2-8=-1 36.0 1.80 750 -500 31.9 2.25 630 31.9 11 -20
11436 10/21/21 07:09-7:57 LH2 3.728 6.5 uA GMN1 ps1=0, ps2-8=-1 36.0 1.80 750 -500 31.9 2.25 630 31.9 11 -20 Production data. Fiber 23 off on GEMs

LD2 Runs

Run Info Beam Settings DAQ BigBite Spectrometer Settings Super BigBite Spectrometer Settings Notes
Num Date Time Target Ebeam (GeV) Ibeam (uA) CODA Config Pre- scales Angle (º) Distance (m) Current (A) B-Field (T) Threshold (mV) SBS Angle (º) SBS Dist (m) Current (A) B-Field (T) HCal Angle (º) HCal Dist (m) Threshold (mV) Notes
11446 10/21/21 10:23-10:30 LD2 3.728 3 uA GMN1 ps1=0, ps2-8=-1 36.0 1.80 750 -500 31.9 2.25 630 31.9 11 -20 Crashed at 600k events on GEM fiber 23. Data is probably still good.
11447 10/21/21 10:35-10:37 LD2 3.728 3 uA GMN1 ps1=0, ps2-8=-1 36.0 1.80 750 -500 31.9 2.25 630 31.9 11 -20
11448 10/21/21 10:38-10:41 LD2 3.728 3 uA GMN1 ps1=0, ps2-8=-1 36.0 1.80 750 -500 31.9 2.25 630 31.9 11 -20
11449 10/21/21 10:45-11:22 LD2 3.728 3 uA GMN1 ps1=0, ps2-8=-1 36.0 1.80 750 -500 31.9 2.25 630 31.9 11 -20 GEM Fiber 23 disabled.
11451 10/21/21 11:26-11:49 LD2 3.728 3 uA GMN1 ps1=0, ps2-8=-1 36.0 1.80 750 -475 31.9 2.25 630 31.9 11 -20 Changed BBCal threshold
11452 10/21/21 11:50-12:45 LD2 3.728 3 uA GMN1 ps1=0, ps2-8=-1 36.0 1.80 750 -500 31.9 2.25 630 31.9 11 -20 Changed BBCal threshold


11456 10/21/21 13:04-14:34 LD2 3.728 1.75 uA GMN1 ps1=0, ps2-8=-1 36.0 1.80 750 -425 31.9 2.25 630 31.9 11 -20 Changed BBCal threshold and beam current to maximize live time

Commissioning Point 1

Commissioning Data with sieve OUT

Note: Calorimeter calibration information

Run Info Beam Settings DAQ BigBite Spectrometer Settings Super BigBite Spectrometer Settings Notes
Num Date Time Target Ebeam (GeV) Ibeam (uA) CODA Config Pre- scales Angle (º) Distance (m) Current (A) B-Field (T) Threshold (mV) SBS Angle (º) SBS Dist (m) Current (A) B-Field (T) HCal Angle (º) HCal Dist (m) Threshold (mV) Notes
11304 10/18/21 00:24-01:22 LH2 1.92 5 uA GMN1 ps1=0, ps2-8=-1 51.0 1.85 750 -275 33.5 2.25 524 34.5 13.5 -20 statistics for SBS at +25%
11305 10/18/21 01:22-01:55 LH2 1.92 5 uA GMN1 ps1=0, ps2-8=-1 51.0 1.85 750 -275 33.5 2.25 524 34.5 13.5 -20 statistics for SBS at +25%
11306 10/18/21 02:23-02:41 LH2 1.92 5 uA GMN1 ps1=0, ps2-8=-1 51.0 1.85 750 -275 33.5 2.25 524 34.5 13.5 -20 statistics for SBS at +25% The beam was down for most of this run, so really this run only has maybe a minute or two of data with beam on.
11307 10/18/21 02:43-02:50 LH2 1.92 5 uA GMN1 ps1=0, ps2-8=-1 51.0 1.85 750 -275 33.5 2.25 524 34.5 13.5 -20 statistics for SBS at +25%
11308 10/18/21 02:52-03:41 LH2 1.92 5 uA GMN1 ps1=0, ps2-8=-1 51.0 1.85 750 -275 33.5 2.25 524 34.5 13.5 -20 statistics for SBS at +25%
11309 10/18/21 03:41-04:16 LH2 1.92 5 uA GMN1 ps1=0, ps2-8=-1 51.0 1.85 750 -275 33.5 2.25 524 34.5 13.5 -20 statistics for SBS at +25%
11310 10/18/21 04:20-05:13 LH2 1.92 5 uA GMN1 ps1=0, ps2-8=-1 51.0 1.85 750 -275 33.5 2.25 524 34.5 13.5 -20 statistics for SBS at +25%
11311 10/18/21 05:17-05:24 LH2 1.92 5 uA GMN1 ps1=0, ps2-8=-1 51.0 1.85 750 -275 33.5 2.25 524 34.5 13.5 -20 statistics for SBS at +25%

Optics Data with sieve IN

Note: These are the important runs for optics analysis:

Run Info Beam Settings DAQ BigBite Spectrometer Settings Super BigBite Spectrometer Settings Notes
Num Date Time Target Ebeam (GeV) Ibeam (uA) CODA Config Pre- scales Angle (º) Distance (m) Current (A) B-Field (T) Threshold (mV) SBS Angle (º) SBS Dist (m) Current (A) B-Field (T) HCal Angle (º) HCal Dist (m) Threshold (mV) Notes
11107 10/14/21 05:06-5:58 5-foil optics 1.92  ?? GMN1 ps1=0, ps2=7 51.0 1.85 0 -275 33.5 2.25 2100 34.5 13.5 -56.5 5-foil optics.
11150 10/14/21 21:51-22:18 Single-foil carbon 1.92 1 GMN1 ps1=0, ps2=7 51.0 1.85 0 -275 33.5 2.25 2100 34.5 13.5 -56.5 zero field optics run for alignment
11153 10/14/21 22:35-23:13 Single-foil carbon 1.92 1 uA GMN1 ps1=0, ps2=7 51.0 1.85 0 -275 33.5 2.25 2100 34.5 13.5 -56.5 zero-field optics run alignment
11154

LH2 with BB Sieve IN

Run Info Beam Settings DAQ BigBite Spectrometer Settings Super BigBite Spectrometer Settings Notes
Num Date Time Target Ebeam (GeV) Ibeam (uA) CODA Config Pre- scales Angle (º) Distance (m) Current (A) B-Field (T) Threshold (mV) SBS Angle (º) SBS Dist (m) Current (A) B-Field (T) HCal Angle (º) HCal Dist (m) Threshold (mV) Notes
11180 10/15/21 03:47-03:54 LH2 1.92 10 GMN1 ps1=0, ps2=7 51.0 1.85 750 -275 33.5 2.25 2100 34.5 13.5 -56.5
11181 10/15/21 03:57-04:05 LH2 1.92 5 GMN1 ps1=0 51.0 1.85 750 -275 33.5 2.25 2100 34.5 13.5 -56.5
11182 10/15/21 04:06-06:05 LH2 1.92 5 GMN1 ps1=0 51.0 1.85 750 -275 33.5 2.25 2100 34.5 13.5 -56.5
11184 10/15/21 06:15-06:21 LH2 1.92 5 GMN1 ps1=0 51.0 1.85 750 -275 33.5 2.25 2100 34.5 13.5 -56.5
11185 10/15/21 06:21-08:21 LH2 1.92 5 GMN1 ps1=0 51.0 1.85 750 -275 33.5 2.25 2100 34.5 13.5 -56.5
11186 10/15/21 08:24-08:44 LH2 1.92 5 GMN1 ps1=0 51.0 1.85 750 -275 33.5 2.25 2100 34.5 13.5 -56.5

LH2 and LD2 with BB Sieve OUT

Run Info Beam Settings DAQ BigBite Spectrometer Settings Super BigBite Spectrometer Settings Notes
Num Date Time Target Ebeam (GeV) Ibeam (uA) CODA Config Pre- scales Angle (º) Distance (m) Current (A) B-Field (T) Threshold (mV) SBS Angle (º) SBS Dist (m) Current (A) B-Field (T) HCal Angle (º) HCal Dist (m) Threshold (mV) Notes
11188 10/15/21 09:39-10:57 None 1.92 No Beam GMN1 ps1=0 51.0 1.85 0 and 750* -275 33.5 2.25 2100 34.5 13.5 -56.5 Cosmics. BB set changed from 0 to 750 A.
11190 10/15/21 10:59-13:58 None 1.92 5* GMN1 ps1=0 51.0 1.85 750 -275 33.5 2.25 2100 34.5 13.5 -56.5 Lots of no beam and short attempt to ramp to 10 uA.
11191 10/15/21 14:06-14:18 C 2mm hole 1.92 1 GMN1 ps1=0 51.0 1.85 750 -275 33.5 2.25 2100 34.5 13.5 -56.5 DAQ problems.
11192 10/15/21 14:21-14:30 C 2mm hole 1.92 1 GMN1 ps1=0 51.0 1.85 750 -275 33.5 2.25 2100 34.5 13.5 -56.5
11193 10/15/21 14:32-14:45 C 2mm hole 1.92 1 GMN1 ps1=0 51.0 1.85 750 -275* 33.5 2.25 2100 34.5 13.5 -56.5 BB threshold changed to -180 mV 4 minutes before run ended.
11195 10/15/21 14:49-14:57 C 2mm hole 1.92 1 GMN1 ps1=0 51.0 1.85 750 -180 33.5 2.25 2100 34.5 13.5 -56.5
11196 10/15/21 14:58-15:03 C 2mm hole 1.92 1 GMN1 ps1=0 51.0 1.85 750 -180 33.5 2.25 2100 34.5 13.5 -56.5
11197 10/15/21 15:16-16:18 LH2 1.92 1 GMN1 ps1=0 51.0 1.85 750 -275 33.5 2.25 2100 34.5 13.5 -56.5
11199 10/15/21 16:31-17:03 LH2 1.92 1 GMN1 ps1=0 51.0 1.85 750 -275 33.5 2.25 2100 34.5 13.5
11201 10/15/21 20:16-20:21 LH2 1.92 1 GMN1 ps1=0 51.0 1.85 750 -251 33.5 2.25 2100 34.5 13.5 -56.5
11202 10/15/21 20:21-20:54 LH2 1.92 1 GMN1 ps1=0 51.0 1.85 750 -251 33.5 2.25 2100 34.5 13.5 -56.5
11203 10/15/21 20:21-20:54 LH2 1.92 1 GMN1 ps1=0 51.0 1.85 750 -251 33.5 2.25 2100 34.5 13.5 -56.5
11204 10/15/21 - 10/16/21 22:50-01:03 LH2 1.92 No Beam GMN1 ps1=0 51.0 1.85 750 -251 33.5 2.25 2100 34.5 13.5 -56.5 Cosmics.
11205 10/16/21 01:20-06:14 LH2 1.92 3, 1* GMN1 ps1=0, ps2=2 51.0 1.85 750 -251 33.5 2.25 2100 34.5 13.5 -56.5 Current was varied from 3 and 1 uA with a lot of no beam.
11207 10/16/21 06:18-07:21 LH2 1.92 1 GMN1 ps1=0, ps2=2 51.0 1.85 750 -251 33.5 2.25 2100 34.5 13.5 -56.5
11208 10/16/21 07:22-08:29 LH2 1.92 1, 2 GMN1 ps1=0, ps2=2 51.0 1.85 750 -251 33.5 2.25 2100 34.5 13.5 -56.5 Current increased from 1 to 2 uA towards run start.
11210 10/16/21 08:35-08:49 LH2 1.92 2 GMN1 ps1=0, ps2=2 51.0 1.85 750 -251 33.5 2.25 2100 34.5 13.5 -56.5
11211 10/16/21 08:50-08:54 LH2 1.92 2 GMN1 ps1=0, ps2=2 51.0 1.85 750 -251 33.5 2.25 2100 34.5 13.5 -56.5
11212 10/16/21 08:54-09:22 LH2 1.92 2 GMN1 ps1=0, ps2=2 51.0 1.85 750 -251 33.5 2.25 2100 34.5 13.5 -56.5
11213 10/16/21 09:24-09:27 LH2 1.92 2 GMN1 ps1=0, ps2=2 51.0 1.85 750 -251 33.5 2.25 2100 34.5 13.5 -56.5
11214 10/16/21 09:30-09:37 LH2 1.92 2 GMN1 ps1=0, ps2=2 51.0 1.85 750 -251 33.5 2.25 2100 34.5 13.5 -56.5
11215 10/16/21 09:55-10:12 LH2 1.92 3 GMN1 ps1=0, ps2=2 51.0 1.85 750 -251 33.5 2.25 2100 34.5 13.5 -56.5
11216 10/16/21 10:18-11:15 LH2 1.92 3 GMN1 ps1=0 51.0 1.85 750 -263 33.5 2.25 2100 34.5 13.5 -56.5
11217 10/16/21 11:31-11:41 LH2 1.92 3 GMN1 ps1=0 51.0 1.85 750 -263 33.5 2.25 2100 34.5 13.5 -56.5 DAQ had to be reset.
11218 10/16/21 11:43-11:48 LH2 1.92 3 GMN1 ps1=0 51.0 1.85 750 -263 33.5 2.25 2100 34.5 13.5 -56.5
11219 10/16/21 11:57-11:59 LH2 1.92 3 GMN1_ Pedestal ps1=0 51.0 1.85 750 -263 33.5 2.25 2100 34.5 13.5 -56.5
11220 10/16/21 12:53-12:55 LH2 1.92 4 GMN1 ps1=0 51.0 1.85 750 -263 33.5 2.25 2100 34.5 13.5 -56.5
11221 10/16/21 12:57-13:01 LH2 1.92 4 GMN1 ps1=0 51.0 1.85 750 -263 33.5 2.25 2100 34.5 13.5 -56.5
11222 10/16/21 13:04-13:12 LH2 1.92 4 GMN1 ps1=0 51.0 1.85 750 -263 33.5 2.25 2100 34.5 13.5 -56.5
11224 10/16/21 13:21-14:23 LH2 1.92 4 GMN1 ps1=0 51.0 1.85 750 -263 33.5 2.25 2100 34.5 13.5 -56.5
11226 10/16/21 14:28-14:41 LH2 1.92 4 GMN1 ps1=0 51.0 1.85 750 -263 33.5 2.25 2100 34.5 13.5 -56.5
11227 10/16/21 14:44-14:46 LH2 1.92 5 GMN1 ps1=0 51.0 1.85 750 -263 33.5 2.25 2100 34.5 13.5 -56.5
11228 10/16/21 14:46-15:06 LH2 1.92 5 GMN1 ps1=0 51.0 1.85 750 -263 33.5 2.25 2100 34.5 13.5 -56.5
11229 10/16/21 15:10-16:23 LH2 1.92 5 GMN1 ps1=0 51.0 1.85 750 -274 33.5 2.25 2100 34.5 13.5 -56.5
11231 10/16/21 16:27-16:31 LH2 1.92 5 GMN1 ps1=0 51.0 1.85 750 -274 33.5 2.25 2100 34.5 13.5 -56.5
11232 10/16/21 16:37-16:39 LH2 1.92 5 GMN1 ps1=0 51.0 1.85 750 -274 33.5 2.25 2100 34.5 13.5 -56.5
11233 10/16/21 16:41-16:44 LH2 1.92 5 GMN1 ps1=0 51.0 1.85 750 -274 33.5 2.25 2100 34.5 13.5 -56.5
11234 10/16/21 16:47-16:49 LH2 1.92 5 GMN1 ps1=0 51.0 1.85 750 -274 33.5 2.25 2100 34.5 13.5 -56.5
11235 10/16/21 16:55-17:13 LH2 1.92 5 GMN1 ps1=0 51.0 1.85 750 -274 33.5 2.25 2100 34.5 13.5 -56.5
11236 10/16/21 18:03-18:09 C 2mm Hole 1.92 1 GMN1 ps1=0 51.0 1.85 750 -175 33.5 2.25 1050 34.5 13.5 -56.5 Spot check for centering.
11237 10/16/21 18:15-18:21 LD2 1.92 No Beam GMN1 ps1=0 51.0 1.85 750 -274 33.5 2.25 1050 34.5 13.5 -56.5 LD2 ion chamber calibration.
11238 10/16/21 18:23-18:28 LD2 1.92 1-10 GMN1 ps1=0 51.0 1.85 750 -274 33.5 2.25 1050 34.5 13.5 -56.5 Ramping current.
11239 10/16/21 18:43-19:16 LD2 1.92 2.5 GMN1 ps1=0 51.0 1.85 750 -274 33.5 2.25 1050 34.5 13.5 -56.5
11240 10/16/21 19:20-19:22 LD2 1.92 2.5 GMN1 ps1=0 51.0 1.85 750 -274 33.5 2.25 1050 34.5 13.5 -56.5
11241 10/16/21 19:27-19:59 LD2 1.92 2.5 GMN1 ps1=0 51.0 1.85 750 -274 33.5 2.25 1050 34.5 13.5 -56.5
11242 10/16/21 20:12-20:35 LH2 1.92 5 GMN1 ps1=0 51.0 1.85 750 -274 33.5 2.25 1050 34.5 13.5 -56.5
11244 10/16/21 20:39-20:53 LH2 1.92 3 GMN1 ps1=0 51.0 1.85 750 -274 33.5 2.25 1050 34.5 13.5 -56.5
11245 10/16/21 20:55-21:19 LH2 1.92 3, 5* GMN1 ps1=0 51.0 1.85 750 -274 33.5 2.25 1050 34.5 13.5 -56.5 Ramped from 3 to 5 uA.
11246 10/16/21 21:22-21:25 LH2 1.92 5 GMN1 ps1=0 51.0 1.85 750 -274 33.5 2.25 1050 34.5 13.5 -56.5
11247 10/16/21 22:25-22:28 LH2 1.92 5 GMN1 ps1=0 51.0 1.85 750 -274 33.5 2.25 1050 34.5 13.5 -56.5
11248 10/16/21 22:29-22:53 LH2 1.92 5 GMN1 ps1=0 51.0 1.85 750 -274 33.5 2.25 1050 34.5 13.5 -56.5
11249 10/16/21 22:54-23:20 LH2 1.92 5 GMN1 ps1=0 51.0 1.85 750 -274 33.5 2.25 1050 34.5 13.5 -56.5
11250 10/16/21 23:21-23:41 LH2 1.92 5 GMN1 ps1=0 51.0 1.85 750 -274 33.5 2.25 1050 34.5 13.5 -56.5
11251 10/16/21 - 10/17/21 23:49-00:52 LH2 1.92 5 GMN1 ps1=0 51.0 1.85 750 -274 33.5 2.25 1050 34.5 13.5 -56.5
11252 10/17/21 00:52-01:54 LH2 1.92 5 GMN1 ps1=0 51.0 1.85 750 -274 33.5 2.25 1050 34.5 13.5 -56.5
11253 10/17/21 02:06-02:10 LD2 1.92 2.5 GMN1 ps1=0 51.0 1.85 750 -274 33.5 2.25 1050 34.5 13.5 -56.5
11254 10/17/21 02:11-03:03 LD2 1.92 2.5 GMN1 ps1=0 51.0 1.85 750 -274 33.5 2.25 1050 34.5 13.5 -56.5, -20* HCal threshold changed from -56.5 -> -20 mV half way.
11255 10/17/21 03:04-03:06 LD2 1.92 2.5 GMN1 ps1=0 51.0 1.85 750 -274 33.5 2.25 1050 34.5 13.5 -20
11256 10/17/21 03:08-03:30 LD2 1.92 2.5 GMN1 ps1=0 51.0 1.85 750 -274 33.5 2.25 1050 34.5 13.5 -20
11257 10/17/21 03:33-03:41 LD2 1.92 2.5 GMN1 ps1=0 51.0 1.85 750 -274 33.5 2.25 1050 34.5 13.5 -20
11258 10/17/21 03:43-04:01 LD2 1.92 2.5 GMN1 ps1=0 51.0 1.85 750 -274 33.5 2.25 1050 34.5 13.5 -20
11259 10/17/21 04:05-04:19 LD2 1.92 2.5 GMN1 ps1=0 51.0 1.85 750 -274 33.5 2.25 1050 34.5 13.5 -20
11260 10/17/21 04:20-04:29 LD2 1.92 2.5 GMN1 ps1=0 51.0 1.85 750 -274 33.5 2.25 1050 34.5 13.5 -20
11261 10/17/21 04:31-04:55 LD2 1.92 2.5 GMN1 ps1=0 51.0 1.85 750 -274 33.5 2.25 1050 34.5 13.5 -20
11262 10/17/21 04:59-05:22 LD2 1.92 2.5 GMN1 ps1=0 51.0 1.85 750 -274 33.5 2.25 1050 34.5 13.5 -20
11263 10/17/21 05:27-06:22 LH2 1.92 5 GMN1 ps1=0 51.0 1.85 750 -274 33.5 2.25 1050 34.5 13.5 -20
11265 10/17/21 06:30-07:21 LH2 1.92 5 GMN1 ps1=0 51.0 1.85 750 -274 33.5 2.25 1050 34.5 13.5 -20
11267 10/17/21 07:27-08:44 LH2 1.92 5 GMN1 ps1=0 51.0 1.85 750 -274 33.5 2.25 1050 34.5 13.5 -20
11268 10/17/21 08:46-09:54 LH2 1.92 5 GMN1 ps1=0 51.0 1.85 750 -274 33.5 2.25 1050 34.5 13.5 -20
11272 10/17/21 10:36-11:51 LD2 1.92 2.5 GMN1 ps1=0 51.0 1.85 750 -274 33.5 2.25 1050 34.5 13.5 -20
11274 10/17/21 12:26-12:29 Carbon hole 1.92 2.5 GMN1 ps1=0 51.0 1.85 750 -174 33.5 2.25 -525 34.5 13.5 -20
11275 10/17/21 12:33-12:36 Carbon hole 1.92 2.5 GMN1 ps1=0 51.0 1.85 750 -174 33.5 2.25 -525 34.5 13.5 -20
11276 10/17/21 12:48-13:20 LH2 1.92 5 GMN1 ps1=0 51.0 1.85 750 -275 33.5 2.25 -525 34.5 13.5 -20
11277 10/17/21 13:25-13:57 LH2 1.92 5 GMN1 ps1=0 51.0 1.85 750 -275 33.5 2.25 -525 34.5 13.5 -20
11279 10/17/21 14:01-14:26 LH2 1.92 5 GMN1 ps1=0 51.0 1.85 750 -275 33.5 2.25 -525 34.5 13.5 -20 DAQ hung, run-11279 was restarted
11280 10/17/21 14:53-14:58 LH2 1.92 5 GMN1 ps1=0 51.0 1.85 750 -275 33.5 2.25 -525 34.5 13.5 -20
11281 10/17/21 15:00-16:24 LH2 1.92 5 GMN1 ps1=0 51.0 1.85 750 -275 33.5 2.25 -525 34.5 13.5 -20 Restarted at 15:04.
11282 10/17/21 16:24-17:11 LH2 1.92 5 GMN1 ps1=0 51.0 1.85 750 -275 33.5 2.25 -525 34.5 13.5 -20
11283 10/17/21 17:14-17:17 LH2 1.92 5 GMN1 ps1=0 51.0 1.85 750 -275 33.5 2.25 -525 34.5 13.5 -20
11284 10/17/21 17:37-17:53 LH2 1.92 5 GMN1 ps1=0 51.0 1.85 375 -275 33.5 2.25 -525 34.5 13.5 -20
11285 10/17/21 18:10-18:13 C 2mm Hole 1.92 1 GMN1 ps1=0 51.0 1.85 750 -175 33.5 2.25 4 34.5 13.5 -20
11286 10/17/21 18:15-18:18 C 2mm Hole 1.92 1 GMN1 ps1=0 51.0 1.85 750 -175 33.5 2.25 4 34.5 13.5 -20
11287 10/17/21 18:28-19:06 LH2 1.92 5 GMN1 ps1=0 51.0 1.85 750 -275 33.5 2.25 4 34.5 13.5 -20
11288 10/17/21 19:21-19:30 LH2 1.92 1 GMN1 ps1=0 51.0 1.85 750 -275 33.5 2.25 4 34.5 13.5 -20 1uA beam on GEMs with new HV.
11289 10/17/21 19:31-19:43 LH2 1.92 1 GMN1 ps1=0 51.0 1.85 750 -275 33.5 2.25 4 34.5 13.5 -20
11290 10/17/21 19:46-19:58 LH2 1.92 2 GMN1 ps1=0 51.0 1.85 750 -275 33.5 2.25 4 34.5 13.5 -20 2uA beam on GEMs on H target.
11291 10/17/21 19:59-20:11 LH2 1.92 3 GMN1 ps1=0 51.0 1.85 750 -275 33.5 2.25 4 34.5 13.5 -20 3uA beam and GEMs on H target.
11292 10/17/21 20:14-20:23 LH2 1.92 4 GMN1 ps1=0 51.0 1.85 750 -275 33.5 2.25 4 34.5 13.5 -20 4uA beam and GEMs on H target.
11296 10/17/21 20:49-21:29 LH2 1.92 5 GMN1 ps1=0 51.0 1.85 750 -275 33.5 2.25 4 34.5 13.5 -20 5uA beam and GEMs on H target. Warnings on F1 TDC, rate dropped
11298 10/17/21 21:36-21:50 LH2 1.92 5 GMN1 ps1=0 51.0 1.85 750 -275 33.5 2.25 4 34.5 13.5 -20

HCal LED Voltage Scan

Run Info DAQ HCal Settings Notes
Num Date Approx #evts CODA Config Pre- scales fADC (Latency, Window) LED Pattern LED Events/Step Pulser Rate Voltage Settings Notes
11319 10/18/21 363k GMN1 ps6=0 (1336,200) 8, 2, 3, 4, 6, 1, 10, 12, 14, 16, 18, 20, 22, 24, 26, 28, 30, 0 10k 1kHz voltages_calibrated_from_run_673.hv
11320 10/18/21 390k GMN1 ps6=0 (1336,200) 8, 2, 3, 4, 6, 1, 10, 12, 14, 16, 18, 20, 22, 24, 26, 28, 30, 0 10k 1kHz voltages_calibrated_from_run_673_CMU+25_JLab+100.hv
11321 10/18/21 377k GMN1 ps6=0 (1336,200) 8, 2, 3, 4, 6, 1, 10, 12, 14, 16, 18, 20, 22, 24, 26, 28, 30, 0 10k 1kHz voltages_calibrated_from_run_673_CMU+50_JLab+200.hv
11322 10/18/21 368k GMN1 ps6=0 (1336,200) 8, 2, 3, 4, 6, 1, 10, 12, 14, 16, 18, 20, 22, 24, 26, 28, 30, 0 10k 1kHz voltages_calibrated_from_run_673_CMU+75_JLab+300.hv
11323 10/18/21 369k GMN1 ps6=0 (1336,200) 8, 2, 3, 4, 6, 1, 10, 12, 14, 16, 18, 20, 22, 24, 26, 28, 30, 0 10k 1kHz voltages_calibrated_from_run_673_CMU+100_JLab+400.hv
11324 10/18/21 390k GMN1 ps6=0 (1336,200) 8, 2, 3, 4, 6, 1, 10, 12, 14, 16, 18, 20, 22, 24, 26, 28, 30, 0 10k 1kHz voltages_calibrated_from_run_673_CMU+125_JLab+500.hv
11325 10/18/21 378k GMN1 ps6=0 (1336,200) 8, 2, 3, 4, 6, 1, 10, 12, 14, 16, 18, 20, 22, 24, 26, 28, 30, 0 10k 1kHz voltages_calibrated_from_run_673_CMU+150_JLab+600.hv
11326 10/18/21 364k GMN1 ps6=0 (1336,200) 8, 2, 3, 4, 6, 1, 10, 12, 14, 16, 18, 20, 22, 24, 26, 28, 30, 0 10k 1kHz voltages_calibrated_from_run_673_CMU+175_JLab+700.hv
11327 10/18/21 360k GMN1 ps6=0 (1336,200) 8, 2, 3, 4, 6, 1, 10, 12, 14, 16, 18, 20, 22, 24, 26, 28, 30, 0 10k 1kHz voltages_calibrated_from_run_673_CMU+200_JLab+800.hv
11328 10/18/21 437k GMN1 ps6=0 (1336,200) 8, 2, 3, 4, 6, 1, 10, 12, 14, 16, 18, 20, 22, 24, 26, 28, 30, 0 10k 1kHz voltages_calibrated_from_run_673_CMU+225_JLab+900.hv
11329 10/18/21 367k GMN1 ps6=0 (1336,200) 8, 2, 3, 4, 6, 1, 10, 12, 14, 16, 18, 20, 22, 24, 26, 28, 30, 0 10k 1kHz voltages_calibrated_from_run_673_CMU+250_JLab+1000.hv
11330 10/18/21 365k GMN1 ps6=0 (1336,200) 8, 2, 3, 4, 6, 1, 10, 12, 14, 16, 18, 20, 22, 24, 26, 28, 30, 0 10k 1kHz voltages_calibrated_from_run_673_CMU+275_JLab+1100.hv
11331 10/18/21 365k GMN1 ps6=0 (1336,200) 8, 2, 3, 4, 6, 1, 10, 12, 14, 16, 18, 20, 22, 24, 26, 28, 30, 0 10k 1kHz voltages_calibrated_from_run_673_CMU+300_JLab+1200.hv